# Enabling Gigantic MIMO Beamforming

### v1

@@ IV-C4 Minimum Computations Per Coherence Block
V Numerical Results
We have proposed MiLAC-aided beamforming as a novel beamforming strategy that offers promising advantages in terms of circuit and computational complexity. In this section, we evaluate the performance and computational complexity of MiLAC-aided beamforming when it is used for R-ZFBF.
Consider a multi-user multiple-input single-output (MISO) system with an $N_{T}$-antenna transmitter and $N_{R}$ single-antenna receivers, as presented in Section IV-A. Here, the transmitter precodes the $N_{R}$ symbols $\mathbf{s}\in\mathbb{C}^{N_{R}\times 1}$ intended to the $N_{R}$ receivers through R-ZFBF. With digital beamforming, the R-ZFBF transmitter is realized by computing $\mathbf{F}_{\text{R-ZFBF}}=\mathbf{H}^{H}(\mathbf{H}\mathbf{H}^{H}+\lambda\mathbf{I}_{N_{R}})^{-1}$ and setting the precoding matrix $\mathbf{W}_{\text{R-ZFBF}}$ such that $[\mathbf{W}_{\text{R-ZFBF}}]_{:,n_{R}}=[\mathbf{F}_{\text{R-ZFBF}}]_{:,n_{R}}/\|[\mathbf{F}_{\text{R-ZFBF}}]_{:,n_{R}}\|_{2}$, for $n_{R}=1,\ldots,N_{R}$, assuming uniform power allocation to the users. For large $N_{R}$ and if the signal-to-noise ratio (SNR) is the same at all receiving antennas, the optimal $\lambda$ is $\lambda=N_{R}\sigma_{\mathbf{n}}^{2}/P_{T}$, where $P_{T}$ is the transmit power, i.e., $\mathbf{s}$ has covariance matrix $\mathbf{C}_{\mathbf{s}}=P_{T}/N_{R}\mathbf{I}_{N_{R}}$, $\sigma_{\mathbf{n}}^{2}$ is the noise power, and we assumed the channel to have unit path gain [16, Chapter 12]. With MiLAC-aided beamforming, the MiLAC can directly compute the R-ZFBF matrix in the analog domain, as explained in Section IV-A. In this case, the columns of $\mathbf{W}_{\text{R-ZFBF}}$ cannot be individually normalized, and the resulting precoding matrix is given by $\mathbf{W}_{\text{R-ZFBF}}=\sqrt{N_{R}}\mathbf{F}_{\text{R-ZFBF}}/\|\mathbf{F}_{\text{R-ZFBF}}\|_{F}$, where the normalization factor has been included to ensure that the transmit power is the same as with digital beamforming.
@@ V Numerical Results
In the considered system, the beamforming matrix is reconfigured per coherence block, where each coherence block includes $\tau$ symbol transmissions. We define the computational complexity as the number of real operations per coherence block. For R-ZFBF at the transmitter, MiLAC-aided beamforming requires $6N_{T}N_{R}$ operations per coherence block, as specified in Remark 2, and no operation is required on a per-symbol basis. Conversely, with digital beamforming, R-ZFBF requires $8(N_{T}N_{R}^{2}+N_{R}^{3}/3)$ operations to design the beamforming matrix at each coherence block (see Remark 2). In addition, the matrix-vector product $\mathbf{W}\mathbf{s}$ is performed on a per-symbol basis, requiring $8N_{T}N_{R}$ real operations per symbol time and making the complexity per coherence block of digital beamforming $8(N_{T}N_{R}^{2}+N_{R}^{3}/3)+8N_{T}N_{R}\tau$.
@@
In Fig. 4, we report the achieved sum rate and required computational complexity for MiLAC-aided and digital R-ZFBF beamforming in the case of different values of $N_{R}$ and $N_{T}$, where $N_{R},N_{T}\in\{256,1024,4096\}$ and $N_{R}\leq N_{T}$, and SNR $P_{T}/\sigma_{\mathbf{n}}^{2}$, where $\text{SNR}\in\{0,10,20\}$ dB, by fixing $\tau=100$ symbols per coherence block. The values are averaged over multiple realizations of independent and identically distributed (i.i.d.) Rayleigh distributed channels with unit path gain. We make the following three observations. First, the sum rate increases with the number of transmit antennas $N_{T}$, the number of receivers $N_{R}$, and the SNR, for both MiLAC-aided and digital beamforming. Besides, the computational complexity increases with the number of transmit antennas $N_{T}$ and receivers $N_{R}$, for both MiLAC-aided and digital beamforming, while is independent of the SNR. Second, with a high number of transmit antennas, i.e., $N_{T}\geq 256$, the performance achieved by a MiLAC computing the R-ZFBF matrix in the analog domain as per Section IV-A is the same as with digital beamforming with uniform power allocation. Thus, we can fully exploit the computational benefits of MiLAC-aided beamforming without sacrificing performance. Third, MiLAC-aided beamforming offers the same performance as digital beamforming for all the considered configurations of $N_{R}$, $N_{T}$, and SNR, with significant gains in computational complexity, up to $7.4\times 10^{3}$ times when $N_{R}=N_{T}=4096$. Under a different perspective, MiLAC-aided beamforming can significantly improve the sum rate over digital beamforming given the same computational complexity. For example, MiLAC-aided beamforming can operate R-ZFBF in a system with $N_{R}=N_{T}=4096$ requiring approximately the same complexity as digital R-ZFBF in a system with $N_{R}=N_{T}=256$.
VI Conclusion
@@
In this paper, we demonstrate the application of MiLAC to wireless communications, highlighting its potential to support communications with thousands of antennas by enabling analog-domain computations and beamforming. We first derive RLS as a special case of the LMMSE estimator that has a practical interest in wireless communications, and illustrate how it can be computed by a MiLAC in the analog domain. Following the capability of a MiLAC to compute RLS, we introduce MiLAC-aided beamforming as a novel beamforming strategy enabled by a MiLAC. Compared to state-of-the-art beamforming strategies, MiLAC-aided beamforming offers unique benefits in terms of hardware and computational complexity. Numerical results support the theoretical insights, showing that MiLAC-aided beamforming can perform R-ZFBF with a remarkable reduction in computational complexity, i.e., up $7.4\times 10^{3}$ times.

### v2

@@ VI-A Performance Evaluation
VI-B Computational Complexity Evaluation
@@
To quantify the benefits of MiLAC-aided beamforming in terms of computational complexity, we consider a MIMO system with $N_{T}$ transmitting antennas and $N_{R}$ receiving antennas, with $N_{T}=N_{R}$, where MiLAC-aided beamforming is used at the transmitter to perform R-ZFBF, ZFBF, and MBF, at the receiver for MMSE, ZF, and MF, and for DFT, as discussed in Section IV-B, IV-C, and IV-D. In the considered system, the beamforming matrix is reconfigured per coherence block, where each coherence block includes $\tau$ symbol transmissions. We define the computational complexity as the number of real operations per coherence block, as analyzed in the following.
  • Zero-forcing. For R-ZFBF and ZFBF at the transmitter, or MMSE and ZF at the receiver, MiLAC-aided beamforming requires $6N_{R}^{2}$ operations per coherence block, as specified in Remarks 2 and 3, and no operation is required on a per-symbol basis. Conversely, with digital beamforming, these strategies require $8(N_{R}^{3}+N_{R}^{3}/3)$ operations to design the beamforming matrix at each coherence block (see Remarks 2 and 3). In addition, a matrix-vector product ($\mathbf{W}\mathbf{s}$ or $\mathbf{G}\mathbf{y}$) is performed on a per-symbol basis, requiring $8N_{R}^{2}$ real operations per symbol time and making the complexity per coherence block of digital beamforming $8(N_{R}^{3}+N_{R}^{3}/3)+8N_{R}^{2}\tau$.
  • Matched filtering. For MBF at the transmitter, or MF at the receiver, MiLAC-aided beamforming requires $4N_{R}^{2}$ operations per coherence block, as given in Remarks 2 and 3, with no operation on a per-symbol basis. Besides, performing these strategies with digital beamforming requires $8N_{R}^{2}$ real operations per symbol time to perform the matrix-vector product $\mathbf{W}\mathbf{s}$ or $\mathbf{G}\mathbf{y}$, while no computation is needed to design the beamforming matrix, making the complexity per coherence block $8N_{R}^{2}\tau$.
  • DFT. Performing the DFT with MiLAC-aided beamforming requires no operation, since the MiLAC is not reconfigured at run-time (see Remark 4). Conversely, computing digitally the DFT requires $34/9N_{R}\log_{2}(N_{R})$ operations per symbol time [37], giving $34/9N_{R}\log_{2}(N_{R})\tau$ operations per coherence block.
In Fig. 7, we show the computational complexity of digital and MiLAC-aided beamforming, as a function of the number of antennas $N_{T}=N_{R}$, by fixing $\tau=100$ symbols per coherence block. For the cases zero-forcing and matched filtering, MiLAC-aided beamforming demonstrates up to $1.5\times 10^{4}$ and $2.0\times 10^{2}$ times lower computational complexity compared to digital beamforming, respectively, when the number of antennas is $N_{R}=8192$. This significant reduction comes because of two reasons. First, MiLAC-aided beamforming does not require the computation of the product $\mathbf{W}\mathbf{s}$, or $\mathbf{G}\mathbf{y}$, on a per-symbol basis, which has significant complexity given the high value of $\tau$ [3]. Second, for zero-forcing, the complexity of MiLAC-aided beamforming scales with $\mathcal{O}(N_{R}^{2})$ instead of $\mathcal{O}(N_{R}^{3})$ since the expensive computation of the beamforming matrix is not required. Besides, in the case of DFT, MiLAC-aided beamforming does not require any computation, saving up to $4.0\times 10^{7}$ operations per coherence block when $N_{R}=8192$. Given this significant reduction in computational complexity, we expect a proportional reduction in central processing unit (CPU) energy consumption and processing time.
VI-C Relationship Between Performance and Computational Complexity
We have shown that MiLAC-aided beamforming can achieve the same performance as digital beamforming in Figs. 5 and 6, and we have quantified its gain in terms of computational complexity in Fig. 7. We now investigate how the achieved performance and the required computational complexity are related in digital and MiLAC-aided beamforming. To this end, we consider multi-user MISO system with an $N_{T}$-antenna transmitter serving $N_{R}$ single-antenna receivers through R-ZFBF. For digital beamforming, we consider the R-ZFBF beamforming matrix to have uniform power allocation, as in Section VI-A. For MiLAC-aided beamforming, we consider the LMMSE-inspired design discussed in Section IV-B, where the computation of the R-ZFBF beamforming matrix is done in the analog domain.
@@ VI-C Relationship Between Performance and Computational Complexity
In Fig. 8, we report the achieved sum rate and required computational complexity for MiLAC-aided and digital R-ZFBF beamforming in the case of different values of $N_{R}$ and $N_{T}$, where $N_{R},N_{T}\in\{256,1024,4096\}$ and $N_{R}\leq N_{T}$, and SNR, where $\text{SNR}\in\{0,10,20\}$ dB. We make the following three observations. First, the sum rate increases with the number of transmit antennas $N_{T}$, the number of receivers $N_{R}$, and the SNR, for both MiLAC-aided and digital beamforming. Besides, the computational complexity increases with the number of transmit antennas $N_{T}$ and receivers $N_{R}$, for both MiLAC-aided and digital beamforming, while is independent of the SNR. Second, with a high number of transmit antennas, i.e., $N_{T}\geq 256$, the performance achieved by a MiLAC computing the R-ZFBF matrix in the analog domain as per Section IV-B is the same as with digital beamforming with uniform power allocation. Thus, we can fully exploit the computational benefits of MiLAC-aided beamforming without sacrificing performance. Third, MiLAC-aided beamforming offers the same performance as digital beamforming for all the considered configurations of $N_{R}$, $N_{T}$, and SNR, with significant gains in computational complexity. Under a different perspective, MiLAC-aided beamforming can significantly improve the sum rate over digital beamforming given the same computational complexity. For example, MiLAC-aided beamforming can operate R-ZFBF in a system with $N_{R}=N_{T}=4096$ requiring approximately the same complexity as digital R-ZFBF in a system with $N_{R}=N_{T}=256$.
VII Conclusion
@@
In Part II of this paper, we demonstrate the application of MiLAC to wireless communications, highlighting its potential to support communications with thousands of antennas by enabling analog-domain computations and beamforming. We first derive five special cases of the LMMSE estimator that have a practical interest in wireless communications, such as the LS and MF, and illustrate how they can be computed by a MiLAC in the analog domain. Following the capability of a MiLAC to compute such special cases, we introduce MiLAC-aided beamforming as a novel beamforming strategy enabled by a MiLAC. Compared to state-of-the-art beamforming strategies, MiLAC-aided beamforming offers the following five unique benefits in terms of performance, hardware complexity, and computational complexity. First, it has the same flexibility as digital beamforming, allowing us to achieve maximum performance. Second, it requires the minimum number of costly RF chains, i.e., equal to the number of transmitted symbols. Third, it can be realized with low-resolution ADCs/DACs since the symbols are precoded fully in the analog domain. Fourth, it does not perform any operation on a per-symbol basis, since it does not require precoding in the digital domain. Fifth, it can offer a significant computational complexity reduction in the case of ZF operations at the transmitter or at the receiver, which can be performed with complexity growing with the square of the antenna number, rather than the cube. Numerical results support the theoretical insights, showing that MiLAC-aided beamforming can offer a remarkable reduction in computational complexity. For example, ZF can be achieved with a complexity reduction of $1.5\times 10^{4}$ times.